Current 10 amp passed througj cuso4 for 32min 10 sec.if 896ml oxy liberated at stp calculate efficiency of cell

Grade:12

2 Answers

Vikas TU
7 years ago
from faraday’s law,
moles x valence factor = i*t/F
moles = i*t/F*n
          = 10*1930/2*96485
         = 0.100 approx.
Moles given at STP = 0.896/22.4 = 0.04
efficiency = (0.1-0.04)/0.1  * 100 => 60%
Pulkit Khandelwal
7 years ago
Nmm thanks buddy.a small mistake valency factor n=4.because 2h2o gives o2 + 4h^+ +4eso the moles=.05 nd the answer is 80%